OpenStudy (anonymous):

This is the most confusing thing I have ever had to do for this class: The sets of numbers 7, 24, 25 and 9, 40, 41 are Pythagorean triples. Use what you know about the Pythagorean Theorem and explain or show why they are Pythagorean triples. Be sure to show your work for each set of triples!

OpenStudy (whpalmer4):

Do you know what a Pythagorean triple is?

OpenStudy (anonymous):

No, I do not.

OpenStudy (whpalmer4):

Okay, that would make this problem pretty difficult :-)

OpenStudy (whpalmer4):

A Pythagorean triple is a set of 3 numbers which could be the lengths of the sides of a right triangle. As the lengths of the sides of a right triangle, they must obey the Pythagorean theorem, which says that the sum of the squares of the lengths of the two shorter sides equals the square of the length of the longest side.

OpenStudy (whpalmer4):

 if you have a right triangle labeled like this, then \[a^2+b^2=c^2\]

OpenStudy (whpalmer4):

So, if you have a set of 3 numbers, if you can substitute them into that formula and get a true statement, they are a Pythagorean triple. The most commonly known one is 3,4,5, because \[3^2 + 4^2 = 5^2\]\[3*3+4*4=5*5\]\[9+16=25\]

OpenStudy (anonymous):

ohh i see

OpenStudy (whpalmer4):

1,2,3 is NOT a Pythagorean triple: \[1^2+2^2=3^2\]\[1*1+2*2=3*3\]\[1+4=9\]\[5=9\]that's not true, so not a Pythagorean triple.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

I don't understand what to put for my question

OpenStudy (anonymous):

9 would be 3^2

OpenStudy (whpalmer4):

Well, one of the triples is 7, 24, 25. You need to show that \[7^2+24^2 = 25^2\]

OpenStudy (anonymous):

so 49 +576 + 625

OpenStudy (whpalmer4):

no, you're missing an = in there

OpenStudy (anonymous):

oh my bad

OpenStudy (whpalmer4):

one of the valuable aspects of math class, even if you never need to use the math you learn again, is that it teaches one to work carefully...

OpenStudy (anonymous):

yes, my bad